# Chypnosis: Stealthy Secret Extraction using Undervolting-based Static Side-channel Attacks

### v2

@@
Abstract.
There is a growing class of static physical side-channel attacks that allow adversaries to extract secrets by probing the persistent state of a circuit. Techniques such as laser logic state imaging (LLSI), impedance analysis (IA), and static power analysis fall into this category. These attacks require that the targeted data remain constant for a specific duration, which often necessitates halting the circuit’s clock. Some methods additionally rely on modulating the chip’s supply voltage to probe the circuit. However, tampering with the clock or voltage is typically assumed to be detectable, as secure chips often deploy sensors that erase sensitive data upon detecting such anomalies. Furthermore, many secure devices use internal clock sources, making external clock control infeasible. In this work, we introduce a novel class of static side-channel attacks, called Chypnosis, that enables adversaries to freeze a chip’s internal clock by inducing a hibernation state via rapid undervolting, and then extracting secrets using static side-channels. We demonstrate that, by rapidly dropping a chip’s voltage below the standard nominal levels, the attacker can bypass the clock and voltage sensors and put the chip in a so-called brownout condition, in which the chip’s transistors stop switching, but volatile memories (e.g., Flip-flops and SRAMs) still retain their data. We test our attack on AMD/Xilinx FPGAs by putting them into hibernation. We show that not only are all clock sources deactivated, but various clock and voltage sensors also fail to detect the tamper event. Afterward, we present the successful recovery of secret bits from a hibernated chip using two static attacks, namely, LLSI and IA. Finally, we discuss potential countermeasures which could be integrated into future designs.
1. Introduction
Figure 1. High-level overview of Chypnosis Attacks
@@ 1. Introduction
Physical Side-Channel Analysis (SCA) attacks and their countermeasures have been extensively studied. These attacks typically exploit the inevitable influence of data transitions during computation on current consumption or voltage drop on a chip. Dynamic side-channel attacks, such as power and electromagnetic analysis, can exploit such data transitions and recover the secret from the chip. Several classes of countermeasures (e.g., masking and hiding) have been proposed and are used in commercial devices to mitigate dynamic SCA attacks. Recently, however, static physical SCA attacks have been gaining attention. In these attacks, adversaries can extract static data stored on memories, such as Flip-Flops (FFs). Examples of such static attacks include static power analysis (Moradi, 2014), Laser Logic State Imaging (LLSI) (Krachenfels et al., 2021a), Impedance Analysis (IA) (Monfared et al., 2023), and Thermal Laser Stimulation(TLS) (Krachenfels et al., 2021b). Some of these techniques, such as LLSI and IA, fall under the category of active sensing or backscatter attacks. In these attacks, the adversary stimulates the chip using external signals (e.g., near-infrared laser beams for LLSI or microwave radiation for IA) and measures the modulated reflections to infer the internal circuit state or memory contents. Due to their active nature, static backscatter attacks often achieve a higher Signal-to-Noise Ratio (SNR) and, in some cases, can extract secrets with a single trace, rendering data randomization techniques like masking ineffective (Krachenfels et al., 2021a; Monfared et al., 2023).
@@
However, such backscatter attacks require some level of tampering with the clock and voltage of the target chip. First, the attacker must freeze the circuit’s state by halting its clock, as more than a clock cycle is required to recover the static data stored in registers (Krachenfels et al., 2021a; Monfared et al., 2023). Second, the adversary must modulate the voltage supplying the chip to produce a detectable modulated reflection during laser or microwave stimulation. Consequently, detecting tampering with the clock and voltage and responding by wiping the sensitive data is the most effective countermeasure against these static attacks. Various clock and voltage sensors have been proposed for this purpose (Dumitru et al., 2025; Farheen et al., 2023). Specialized sensors targeting specific threat vectors, such as voltage glitching (Zick et al., 2013; Provelengios et al., 2021) and laser probing (Khalaj Monfared et al., 2024b; Zhang et al., 2023; Tajik et al., 2017a), can serve as effective countermeasures against backscatter attacks. Nonetheless, assuming that an attacker can access and manipulate the system clock is often unrealistic in real-world scenarios. Many secure Integrated Circuits (ICs) rely on internal clock sources for cryptographic operations, making external clock control extremely difficult for an adversary.
@@
Driven by these observations, we raise the following research questions: (1) Is it possible to halt the system’s clock without tampering with its source? (2) Can static side-channel attacks be performed successfully without triggering clock and voltage sensors?
@@
Contribution. In this work, we positively answer the above questions. We introduce Chypnosis attacks, in which an adversary places a chip in hibernation, bypasses clock and voltage sensors, and recovers secret data using backscatter techniques. Our attack exploits the observation that rapidly lowering the supply voltage below nominal thresholds induces a brownout condition, where logic components (e.g., state machines and clock buffers) stop switching, but volatile memory elements such as SRAM and flip-flops continue to retain data. Fig. 1 presents an abstract overview of our attack.
@@
To validate this attack, we perform Chypnosis on AMD/Xilinx FPGAs fabricated using a 28 nm process. First, we conduct extensive experiments to determine the voltage thresholds required to induce hibernation at various clock frequencies. Next, we demonstrate that entering this brownout condition effectively halts the clock and suppresses switching activity without requiring direct control over the clock source. We confirm this behavior through electrical and optical measurements.
We further show that this switching freeze disables the response circuitry of recently proposed clock sensors on FPGAs (Dumitru et al., 2025; Farheen et al., 2023), which were designed to defend against static SCA attacks. Additionally, we demonstrate how our rapid voltage drop deactivates the response mechanism of existing voltage sensors on FPGAs (Peterson, 2017), which are supposed to detect voltage tampering attacks.
Finally, we show that even in the brownout state, it is possible to delicately modulate the supply voltage, without crashing or waking the system, to successfully perform LLSI and IA attacks and recover secret data in a single trace, see Fig. 1. We conclude by discussing potential countermeasures, which, unfortunately, often require external system modifications or are incompatible with legacy hardware.
2. Technical Background
Figure 2. High-level overview of LLSI setup
2.1. Laser Logic State Imaging (LLSI)
Laser Logic State Imaging (LLSI) (Niu et al., 2014) is a subset of optical probing techniques originally developed for failure analysis of ICs, which have recently gained attention for the potential of side-channel attacks. In optical probing, a near-infrared laser is focused on the transistors from the backside of an IC, where its reflection becomes modulated by the gate or drain of a transistor during switching activity. This modulated reflection can be analyzed in two main ways. In the first, known as Laser Voltage Probing (LVP) or Electro-Optical Probing (EOP), the attacker repeatedly samples the reflection at a single point to reconstruct a waveform of the processed data. In the second, known as Laser Voltage Imaging (LVI) or Electro-Optical Frequency Mapping (EOFM), the laser is scanned across a region of interest while a spectrum analyzer filters out modulations at a specific frequency. In a typical LVI setup, the objective is to identify transistors switching at a known frequency and generate a 2D activity map highlighting regions operating at that frequency. Combined with LVP, these techniques can reveal internal signal activity, provided that the signals are not static (Tajik et al., 2017b; Chef et al., 2022).
LLSI builds upon the LVI technique by enabling the probing or imaging of static signals. By modulating the power supply at a known frequency, as illustrated in Fig. 2, the voltage at the transistors that are in ON states will also be modulated, generating a measurable LVI signal. In contrast, transistors in the OFF state do not produce a significant signal. As a result, the logic state of a memory cell can be inferred based on the observed LVI activity. Fig. 2 presents a simplified example of an LLSI image of an SRAM cell composed of two cross-coupled inverters. LLSI has been successfully used to extract data from registers in FPGAs (Krachenfels et al., 2021a, b) and SRAM cells in microcontrollers(Kiyan et al., 2018; Chef et al., 2021).
2.2. Impedance Analysis (IA)
Impedance Analysis (IA) (Monfared et al., 2023) is a non-invasive side-channel attack that can recover secret data (like cryptographic keys) by measuring changes in the impedance of an IC’s power delivery network (PDN). The key idea is that the temporary contents of registers and their corresponding wiring influence the IC’s physical characteristics (i.e., impedance), which leads to changes in how electrical signals with various frequencies reflect or transmit through the IC’s PDN. To measure such reflections and transmissions, the attacker uses a Vector Network Analyzer (VNA) to inject high-frequency sine waves into the IC’s power rails and then captures the so called scattering parameters. Different regions of an IC respond differently at various frequencies (Mosavirik et al., 2023a), and thus, by sweeping across a range of frequencies, the attacker can essentially probe multiple areas of the chip simultaneously without needing physical access to specific wires or locations.
Fig. 3 illustrates a high-level overview of an impedance attack. The attack process is conceptually similar to channel estimation in wireless communications, where reference radio frequency (RF) signals are transmitted through a channel and the received signals are analyzed. In the case of impedance attacks, the ”channel” corresponds to the power delivery network (PDN) of the target system, while the transmitter and receiver are the ports of a VNA. RF waves are injected into the PDN at specific frequencies and amplitudes, and the responses are captured at the receiver with amplitude and phase modulations introduced by the circuit’s internal state. By analyzing these modulated parameters, an attacker can extract secret information. As with other side-channel attacks, both profiling and non-profiling attacks (Chari et al., 2002) can be applied to impedance analysis by collecting and processing impedance traces. Prior work has demonstrated the effectiveness of this approach for attacking both protected and unprotected cryptographic implementations (Monfared et al., 2023), as well as for reverse engineering purposes (Awal and Rahman, 2023).
2.3. Conventional Countermeasures
Countermeasures such as data and instruction randomization, commonly implemented through masking or hiding, are widely used to mitigate dynamic side-channel attacks by preventing the attacker from integrating measurements over multiple clock cycles. However, these techniques become ineffective when an attacker can freeze the IC’s state and simultaneously probe multiple registers in a single-trace attack, such as with LLSI or IA. To successfully launch an SCA attack, specific conditions must be met. Several countermeasures have been proposed in the literature to prevent these conditions from being satisfied, thus mitigating SCA threats. Here, we review some of these countermeasures.
Figure 3. High-level illustration of impedance attack
2.3.1. Detection-based Countermeasures
@@
Clock Sensors. The first and foremost requirement to launch a static SCA attack is to stop the system clock. Hence, if we have a sensor that detects that the clock has been halted for a while, it can trigger a response, such as wiping sensitive data, before an adversary can recover it. Some secure IC families, such as microcontrollers used in smartcards, are equipped with internal clock sensors capable of detecting anomalies in clock behavior. However, the specific design details of these sensors are proprietary. FPGAs also contain internal circuitry, such as phase-locked loops (PLLs), that can detect irregularities in the clock waveform, but they must be explicitly configured by the user as a security sensor. Commercial solutions, such as the AMD/Xilinx SecMon IP core, offer built-in clock monitoring capabilities, though such features are typically available only to defense and aerospace customers.
There have been a few attempts in the literature to design clock freeze detection sensors for FPGAs. For instance, Farheen et al.(Farheen et al., 2023) proposed using internal clock sources (e.g., a ring oscillator) to monitor the integrity of external clocks. More recently, Dumitru et al.(Dumitru et al., 2025) introduced two sensor variants that detect clock freezing without relying on any internal clock sources.
@@
Voltage Sensors. Similar to clock sensors, many secure ICs are equipped with voltage sensors to detect voltage tampering. On FPGAs, analog-to-digital converters (ADCs), such as the XADC available in AMD/Xilinx FPGAs (Xilinx Inc., 2025), provide built-in voltage sensing capabilities. These sensors can monitor both internal and external voltages, converting analog signals into digital values that can be processed by user-defined digital logic on the FPGA. In addition to built-in voltage sensors, FPGA users can also configure their own delay-based ADCs, such as ring oscillators (ROs) (Zick and Hayes, 2012) and time-to-digital converters (TDCs) (Zick et al., 2013), on the FPGA.
@@
Laser Sensors. For certain backscatter-based attacks, such as LLSI, sensors capable of directly detecting incident laser beams have also been investigated. Similarly to voltage tampering detection, ADCs have demonstrated sensitivity to localized temperature increases caused by laser irradiation. As a result, ROs (Tajik et al., 2017a) and TDCs (Khalaj Monfared et al., 2024b) have been utilized to detect laser probing attacks on FPGAs.
2.3.2. Response-based Countermeasures
@@
An often overlooked aspect of countermeasures is the system’s response after such powerful attacks are detected. Although the conventional assumption is that sensitive data can simply be zeroized upon detection, this may not be suitable in many real-world scenarios. First, zeroization itself can cause dynamic side-channel leakage, such as through power side-channels that reveal Hamming weights. To mitigate this, operations such as masked clear are employed, in which sensitive register contents are wiped by overwriting them with random values (Dumitru et al., 2025). Second, sometimes a continued system operation is desired, which could become impossible after zeroization. In such cases, schemes such as Moving Target Defense (MTD), can be employed to mitigate the threat without any interruption in the system’s operation. For instance, randomizing the placement and routing of a circuit on an FPGA via partial reconfiguration has been shown to be effective against LLSI and IA attacks (Khalaj Monfared et al., 2024b, a).
2.4. Brownout Condition
@@
When the source voltage of a transistor is above a certain threshold, the transistor effectively functions like a switch responding to the changes that come from gate voltage. Reducing the supply voltage, commonly referred to as voltage scaling, has been widely used to improve the energy efficiency of microprocessors. The lower bound for voltage scaling is typically constrained to around half of the nominal operating voltage (Haider and Nazhandali, 2008). However, it has been shown that standard CMOS logic gates can even operate well below the threshold voltage. Based on these observations, prototype designs have shown that by carefully replacing analog-like components with standard digital switching elements, it is possible to push voltage scaling into the subthreshold region and extend the traditional limits of low-voltage operation (Sze et al., 2006; Zhai et al., 2009).
@@
On commercial FPGAs, however, subthreshold operation is hard to achieve due to the high energy consumption of conventional FPGA interconnects. While there have been multiple research proposals to enhance the performance of subthreshold FPGAs by optimizing interconnect drivers and operating them in the near-threshold voltage region (Pable and Hasan, 2011; Qi et al., 2016; Grossmann et al., 2012), to the best of our knowledge, they have never been realized on commercial FPGAs. Thus, as the supply voltage drops in FPGAs, a brownout condition occurs where the transistors cannot drive the capacitive loads at the gates of other transistors. Consequently, clock buffers and PLL circuits will also stop functioning, and the clock signal distribution will halt. Meanwhile, memory cells, such as SRAM cells and flip-flops (FFs), may fully retain their data, as their Data Retention Voltage (DRV) (Calhoun and Chandrakasan, 2006; Holcomb et al., 2012) is typically lower than the logic operating threshold. For instance, in the case of SRAM-based FPGAs, a brownout condition leads to the stoppage of all switching activities inside configuration logic blocks (CLBs) and switch boxes, but the SRAMs and FFs retain the FPGA configuration and user data. If the voltage drops further, memory cells will lose their content, and thus, the FPGA will crash and require reconfiguration. Therefore, a narrow subregion within the subthreshold operating range, above the data retention voltage (DRV), exists where the FPGA enters a ”hibernation” state. In this state, switching activity effectively stops, yet all data remains retained. Furthermore, an FPGA can wake up from hibernation by raising the voltage back above the threshold voltage and resume operation as usual.
An attempt was made to read the SRAM content in brownout conditions using laser stimulation (Nedospasov et al., 2013). However, this attack triggers the watchdog timers and brownout event detectors and thus zeroizes all clocked memories with reset lines (e.g., FFs) except SRAMs. Hence, by encrypting the SRAM, one can ensure that no useful data can be extracted. However, our proposed attack (see Sect. 6) bypasses the sensors, prevents data erasure from FFs, and recovers data from them.
3. Threat Model

### v5

# Chypnosis: Undervolting-based Static Side-channel Attacks

@@
Abstract
Static side-channel analysis attacks, which rely on a stopped clock to extract sensitive information, pose a growing threat to embedded systems’ security. To protect against such attacks, several proposed defenses aim to detect unexpected variations in the clock signal and clear sensitive states. In this work, we present Chypnosis, an undervolting attack technique that indirectly stops the target circuit clock, while retaining stored data. Crucially, Chypnosis also blocks the state clearing stage of prior defenses, allowing recovery of secret information even in their presence. However, basic undervolting is not sufficient in the presence of voltage sensors designed to handle fault injection via voltage tampering. To overcome such defenses, we observe that rapidly dropping the supply voltage can disable the response mechanism of voltage sensor systems. We implement Chypnosis on various FPGAs, demonstrating the successful bypass of their sensors, both in the form of soft and hard intellectual property (IP) cores. To highlight the real-world applicability of Chypnosis, we show that the alert handler of the OpenTitan root-of-trust, responsible for providing hardware responses to threats, can be bypassed. Furthermore, we demonstrate that by combining Chypnosis with static side-channel analysis techniques, namely laser logic state imaging (LLSI) and impedance analysis (IA), we can extract sensitive information from a side-channel protected cryptographic module used in OpenTitan, even in the presence of established clock and voltage sensors. Finally, we propose and implement an improvement to an established FPGA-compatible clock detection countermeasure, and we validate its resilience against Chypnosis.
footnotetext: This article has been accepted for publication in IEEE Symposium on Security and Privacy 2026.
1 Introduction
@@
Physical side-channel attacks (SCA) can undermine the security of cryptographic implementations on integrated circuits (ICs). These attacks typically exploit the inevitable influence of data transitions during computation on current consumption or voltage drop on a chip. Dynamic side-channel attacks, such as power [23] and electromagnetic analysis [2], can exploit such data transitions and recover the secret from the chip. Recently, however, static physical SCA attacks have been gaining attention, in which adversaries can extract static data stored in memories, such as Flip-Flops (FFs). Examples of such static attacks include static power analysis [35], Laser Logic State Imaging (LLSI) [24], Impedance Analysis (IA) [31], and Thermal Laser Stimulation (TLS) [25].
Such static attacks require some level of tampering with the clock and voltage of the target chip. First, the attacker must freeze the circuit’s state by halting its clock, because the time required for recovering the static data stored in registers is significantly longer than the clock period [24, 31]. Second, in a subclass of these static attacks known as backscatter attacks (e.g., LLSI and IA), the adversary must modulate the voltage supplying the chip to produce a detectable modulated reflection during laser or microwave stimulation. In these attacks, the adversary stimulates the chip using external signals (e.g., near-infrared laser beams for LLSI or microwave radiation for IA) and measures the modulated reflections to infer the internal circuit state or memory contents. Due to their active nature, static backscatter attacks often achieve a higher Signal-to-Noise Ratio (SNR) and, in some cases, can extract secrets with a single trace, rendering data randomization techniques such as masking ineffective [24, 31].
Figure 1: High-level overview of Chypnosis Attacks
@@
Consequently, detecting clock or voltage tampering and then responding by wiping the sensitive data is a promising countermeasure against these static attacks. Various clock and voltage sensors, both as soft intellectual property (IP) cores [13, 14] and hard IP cores [48, 27], have been developed for this purpose. Moreover, specialized sensors targeting specific threat vectors, such as voltage glitching [63, 49] and laser probing [33, 61, 53], can also serve as effective countermeasures against backscatter attacks. Additionally, the assumption that an attacker can access and manipulate the system clock is often unrealistic in real-world scenarios. Many secure ICs rely on internal clock sources [27] for cryptographic operations, making external clock control extremely difficult for an adversary.
@@
Driven by these observations, in this paper, we ask the following questions: (1) Is it possible to halt the system’s clock without tampering with its source? (2) Can we perform static side-channel attacks without triggering clock and voltage sensors?
1.1 Our Contribution
@@
In this work, we positively answer both questions. We introduce Chypnosis attacks, in which an adversary bypasses clock and voltage sensors, places a chip in hibernation while retaining its data. In this condition, the adversary can perform a static side-channel attack to recover the retained data. Our attack exploits the observation that rapidly lowering the supply voltage below nominal thresholds induces a brownout condition, where logic components (e.g., state machines and clock buffers) cease switching, but volatile memory elements, such as SRAM and flip-flops, continue to retain data. Figure 1 presents an abstract overview of our attack.
@@
We perform Chypnosis on both SRAM-based and Flash-based FPGAs fabricated using 28 nm processes. First, we conduct extensive experiments to determine the voltage thresholds required to induce hibernation at various clock frequencies. Next, we demonstrate that entering this brownout condition effectively halts the clock and suppresses switching activity without requiring direct control over the clock source. We confirm this behavior through electrical and optical measurements. We further show that a switching freeze caused by a rapid voltage drop disables the response circuitry of clock, voltage, and brownout detection (BOD) sensors on FPGAs, irrespective of being implemented as soft IPs [13, 14] or hard IPs [27]. Hence, the sensitive data is not wiped and remains on the chip.
To demonstrate Chypnosis’s effectiveness in practical scenarios, we successfully attack the FPGA implementation of OpenTitan Earl Grey and bypass its alert handler [42], which is responsible for providing hardware responses to threats. Finally, we demonstrate that even in the brownout state, it is possible to delicately modulate the supply voltage without crashing or waking the system, thereby successfully performing LLSI and IA attacks. We demonstrate the extraction of secret data from a side-channel-protected symmetric cryptographic module, used in OpenTitan, in a single trace. We conclude by proposing a circuit-based FPGA-compatible countermeasure, which successfully mitigates Chypnosis.
2 Technical Background
2.1 Conventional Static SCA Countermeasures
2.1.1 Detection-based Countermeasures
@@
Clock Sensors.  The primary requirement for launching a static SCA attack is to stop the system clock. Hence, if we have a sensor that detects that the clock has been halted for a while, it can trigger a response, such as wiping sensitive data, before an adversary can recover it. Some secure IC families are equipped with internal clock sensors capable of detecting anomalies in clock behavior [27]. However, the specific design details of these sensors are proprietary. FPGAs also contain internal circuitry, such as phase-locked loops (PLLs), that can detect irregularities in the clock waveform; however, they must be explicitly configured by the user to function as a security sensor. Commercial solutions, such as the AMD/Xilinx Security Monitor IP core [58], offer built-in clock monitoring capabilities; however, these features are typically available only to specific customers. Several attempts have been made in the literature to design clock freeze detection sensors in the form of soft IPs for FPGAs. For instance, Farheen et al.[14] proposed using internal clock oscillators to monitor the integrity of external clocks. More recently, Dumitru et al.[13] introduced two sensor variants that detect clock freezing without relying on any internal clock sources.
@@
Voltage Sensors.  Similar to clock sensors, many secure ICs are equipped with voltage sensors to detect voltage tampering [27]. On FPGAs, analog-to-digital converters (ADCs), such as the XADC available in AMD/Xilinx FPGAs [59], provide built-in voltage sensing capabilities. These sensors can monitor both internal and external voltages, converting analog signals into digital values that can be processed by user-defined digital logic on the FPGA. In addition to built-in voltage sensors, FPGA users can also configure their own delay-based ADCs, such as ring oscillators (ROs) [62] and time-to-digital converters (TDCs) [63], on the FPGA.
@@
Laser Sensors.  For specific backscatter-based attacks, such as LLSI, sensors capable of directly detecting incident laser beams have also been investigated. Similarly to voltage tampering detection, ADCs have demonstrated sensitivity to localized temperature increases caused by laser irradiation. As a result, ROs [53] and TDCs [33] have been utilized to detect laser probing attacks on FPGAs.
2.1.2 Response-based Countermeasures
@@
An often overlooked aspect of countermeasures is the system’s response after such powerful attacks are detected. Although the conventional assumption is that sensitive data can simply be zeroized upon detection, this may not be suitable in many real-world scenarios. First, zeroization itself can cause dynamic side-channel leakage [13], such as through power side-channels that reveal Hamming weights. To mitigate this, operations such as masked clear are employed, in which sensitive register contents are overwritten with random values [13]. Second, sometimes continued system operation is desired, which could become impossible after zeroization. In such cases, schemes such as Moving Target Defense (MTD) can be employed to mitigate the threat without interrupting the system’s operation. For instance, randomizing the placement and routing of a circuit on an FPGA via partial reconfiguration has been shown to be effective against LLSI and IA attacks [33, 32].
2.2 Brownout Condition
@@
When the source voltage of a transistor exceeds a certain threshold, the transistor effectively functions like a switch, responding to changes in the gate voltage. Reducing the supply voltage, commonly referred to as voltage scaling, has been widely used to improve the energy efficiency of microprocessors. The lower bound for voltage scaling is typically constrained to around half of the nominal operating voltage [17]. However, it has been demonstrated that standard CMOS logic gates can operate effectively even below the threshold voltage. Based on these observations, prototype designs have shown that by carefully replacing analog-like components with standard digital switching elements, it is possible to push voltage scaling into the subthreshold region and extend the traditional limits of low-voltage operation [52, 60]. On commercial FPGAs, however, subthreshold operation is hard to achieve due to the high energy consumption of conventional FPGA interconnects. While multiple research proposals have been proposed to enhance the performance of subthreshold FPGAs by optimizing interconnect drivers and operating them in the near-threshold voltage region [47, 50, 16], they have yet to be realized on commercial FPGAs.
@@
As the supply voltage drops in FPGAs, a brownout condition occurs where the transistors cannot drive the capacitive loads at the gates of other transistors. Consequently, clock buffers and PLL circuits will also cease to function, and the clock signal distribution will halt. Meanwhile, memory cells, such as SRAM cells and flip-flops (FFs), may fully retain their data, as their Data Retention Voltage (DRV) [8, 20] is typically lower than the logic operating threshold. For instance, in the case of SRAM-based FPGAs, a brownout condition causes the stoppage of all switching activities within configuration logic blocks (CLBs) and switch boxes. At the same time, the SRAMs and flip-flops (FFs) retain the FPGA configuration and user data. If the voltage drops further, memory cells will lose their content, and thus, the FPGA will crash and require reconfiguration. Therefore, a narrow subregion within the subthreshold operating range, above the data retention voltage (DRV), exists where the FPGA enters a “hibernation” state. Furthermore, an FPGA can wake up from hibernation by raising the voltage back above the threshold voltage and resume operation as usual.
Figure 2: High-level overview of LLSI attack
2.3 Laser Logic State Imaging (LLSI)
Laser Logic State Imaging (LLSI) [41] is a static side-channel attack method. LLSI is a subset of optical probing attacks, in which a near-infrared laser is focused on the transistors from the backside of an IC, and its reflection becomes modulated by the gate or drain of a transistor during switching activity. This modulated reflection can be analyzed in two main ways. In the first, known as Laser Voltage Probing (LVP), the attacker repeatedly samples the reflection at a single point to reconstruct a waveform of the processed data. In the second, known as Laser Voltage Imaging (LVI), the laser is scanned across a region of interest while a spectrum analyzer filters out modulations at a specific frequency. In a typical LVI setup, the objective is to identify transistors switching at a given frequency and generate a 2D activity map highlighting regions operating at that frequency. Combined with LVP, these techniques can reveal internal signal activity, provided that the signals are not static [54, 11].
LLSI builds upon the LVI technique by enabling the probing or imaging of static signals. By modulating the power supply at a known frequency, as illustrated in Figure 2, the voltage at the transistors that are in ON states will also be modulated, generating a measurable LVI signal. In contrast, transistors in the OFF state do not produce a significant signal. As a result, the logic state of a memory cell can be inferred based on the observed LVI activity. Figure 2 presents a simplified example of an LLSI image of an SRAM cell composed of two cross-coupled inverters. LLSI has been successfully used to extract data from registers in FPGAs [24, 25] and SRAM cells in microcontrollers [22, 10].
2.4 Impedance Analysis (IA)
Impedance Analysis (IA) [31] is another static side-channel attack that can recover secret data by measuring changes in the impedance of an IC’s power delivery network (PDN). The key idea is that the temporary contents of registers and their corresponding wiring influence the IC’s PDN impedance, which leads to changes in how electrical signals with various frequencies reflect or transmit through the IC’s PDN. To measure such reflections and transmissions, the attacker uses a Vector Network Analyzer (VNA) to inject high-frequency sine waves into the IC’s power rails and then captures the so-called scattering parameters. Different regions of an IC respond differently at various frequencies [36], and thus, by sweeping across a range of frequencies, the attacker can essentially probe multiple localized areas of the chip simultaneously without needing physical access to specific wires or locations.
Figure 3 illustrates a high-level overview of an impedance attack. The attack process is conceptually similar to channel estimation in wireless communications, where reference radio frequency (RF) signals are transmitted through a channel and the received signals are analyzed. In the case of impedance attacks, the “channel” corresponds to the PDN of the target system, while the transmitter and receiver are the ports of a VNA. RF waves are injected into the PDN at specific frequencies and amplitudes, and the responses are captured at the receiver with amplitude and phase modulations introduced by the circuit’s internal state. By analyzing these modulated parameters, an attacker can extract secret information. Prior work has demonstrated the effectiveness of this approach for attacking both protected and unprotected cryptographic implementations [31], as well as for reverse engineering purposes [6].
Figure 3: High-level illustration of impedance attack
3 Threat Model